# Gamma-ray burst Jet Progenitor: long vs short bursts by angular momentum and mass ratio

### v1

@@ 6.3 Photon-pair sphere/jet energy and spectrum peak
We have to point out that the photon-pair sphere/jet energetic properties (6.4,6.5) and (6.10,6.12) explicitly depend on the initial and final core central densities $\rho_{c}^{\rm in}$ and $\rho_{c}^{\rm max}$. They characterize the gravitation gain and conversion to baryon kinetic energy and photon-pair thermal energy in gravitational collapses. These characteristics should also be held for gravitational collapses of non-spherical cores or coalescence of binary systems by introducing effective initial and final central densities $(\rho_{c}^{\rm in})_{\rm eff}$ and $(\rho_{c}^{\rm max})_{\rm eff}$ of the system.
@@
By simplifying out $(\rho_{c}^{\rm in})$ and $(\rho_{c}^{\rm max})$ among Eqs. (6.4,6.5) and (6.10,6.12), we obtained theoretical correlations among the spectral peak $E_{p}$, total photon-pair energy $E_{\gamma}$ and peak luminosity $L_{\gamma}\propto cR_{\gamma}^{2}\langle T_{\gamma}\rangle^{4}\propto cE_{\gamma}/R_{\gamma}$,
@@
with the unique positive parameter $\chi=\frac{\gamma-1}{4\gamma-5}\sim{\mathcal{O}}(1)$  [^3] It is an inverse of the original $\chi-$definition [17]. Although many intermediate processes have not been considered in this simplified model, the scaling laws (6.14) of GRBs’ progenitors give an insight into the Amati and Yonetoku empirical relations, as well as the $L_{\gamma}\propto E_{\gamma}^{1+\chi}$ relation in a unified framework.
$J/M$ $\theta_{\rm jet}$ $E^{j}_{p}/E_{p}$ $E^{j}_{\gamma}/E_{\gamma}$ $L^{j}_{\gamma}/L_{\gamma}$ $R^{j}_{\gamma}/R_{\gamma}$ $N^{j}_{\gamma}/N_{\gamma}$ $V^{j}_{\gamma}/V_{\gamma}$
@@
Table 1: For selected values of the rotation parameter ${\mathcal{A}}^{2}a^{2}\Omega$, i.e., the ratio $J/M$ (5.31), we tabulate the jet angle $\theta_{\rm jet}$ (6.6) and the ratios of the spectral peak $E^{j}_{p}/E_{p}$, total energy $E^{j}_{\gamma}/E_{\gamma}$, maximal luminosity $L^{j}_{\gamma}/L_{\gamma}\propto(E^{j}_{\gamma}/E_{\gamma})(R_{\gamma}/R^{j}_{\gamma})$, minimal size $R^{j}_{\gamma}/R_{\gamma}$, total particle number $N^{j}_{\gamma}/N_{\gamma}$ and volume $V^{j}_{\gamma}/V_{\gamma}$ between the photon-pair sphere and jet cases for comparisons. The photon-pair jet (with superscript $j$) and sphere quantities are given by Eqs. (6.4,6.11,6.12). Their ratios do not explicitly depend on the initial (final) core centre density $\rho_{c}^{\rm in}$ ($\rho_{c}^{\rm fi}$) of gravitational collapses.
7 Long vs short GRBs by angular momentum and mass ratio
7.1 Trends of jet features varying with $J/M$
@@
7.2 A separatrix in angular momentum $J/M$
The photon-pair (approximate) sphere case represents one distinct class of the GRBs progenitors with small angular momenta ($J/M\gtrsim 0$), which are more likely for the cases of massive core collapses, relating to long GRBs. The photon-pair jet case with nontrivial angular momenta represents another distinct class of the GRBs progenitors, which are more likely for the cases of binary coalescence, relating to short GRBs. We will come back to this point later. In the cases of binary coalescence, systems have larger angular momenta $J/M$, where the core mass $M$ should be an effective value of total system mass, the central density $\rho_{c}$ (5.11) and its initial value (5.27) should be effective one. Since the separation between long and short GRBs classes is distinct, there must be a separatrix in angular momentum $J/M$. It occurs when the centrifugal potential $\phi_{c}$ is too large to be balanced by the gravitational potential $\phi$ and the material binding energy $\lambda$. Using the equilibrium equation (5.9) and gravitational potential $\phi$ balanced by the heat function $h$, we give a separatrix,
$\phi_{c}\approx\frac{(v_{s}^{c})^{2}}{\gamma-1}\frac{\lambda}{6}r^{2}\quad\Rightarrow\quad J/M\propto({\mathcal{A}}^{2}a^{2}\Omega)_{\rm sepa}\sim\frac{\lambda}{3(\gamma-1)}\approx\lambda$ (7.1)
in this simplified model. When the repulsive rotation energy is larger than the material binding energy, namely $J/M>\lambda$, they are for the cases of binary coalescence. On the contrary, $J/M<\lambda$ are for the cases of massive core collapses. The real situation is much more complex.
7.3 Long vs short GRBs
The following qualitative observations and comments are in order from the trends of increasing or decreasing quantities as increasing angular momentum $J/M$ in Table 1.
@@ 7.3 Long vs short GRBs
    Due to the angular momentum and repulsive centrifugal potential increase, the gravitational energy gain becomes smaller. The ratios $E_{\gamma}^{j}/E_{\gamma}$ and $N_{\gamma}^{j}/N_{\gamma}$ are smaller than one but still in the order of unity, because most photons and pairs are produced at the centre ($z=0,y=0$) of the sphere and jet, see Fig. 3.
@@
    The jet spectral peak $E_{p}^{j}/E_{p}$ significantly increases up to twice larger than the spherical one, because the photon-pair energy density increases as the jet volume $V^{j}$ decreases. As a result, the spectral peak $E^{j}_{p}$ increases and total energy $E^{j}_{\gamma}$ decreases as angular momentum increases. It gives a possible explanation for two distinct Amati relations $E_{p}\propto E_{\gamma}^{\chi}$ (6.14) respectively for long and short bursts’ data scattering on the $E_{p}$-$E_{\gamma}$ plane, see Fig 7 (right) of the reference [25], where two distinct Amati relations are observed for long and short bursts.
@@
    The spectral peak $E^{j}_{p}$ and the peak luminosity $L^{j}_{\gamma}$ increase as angular momentum increases. It implies the Yonetoku relation $E_{p}\propto L_{\gamma}^{\chi/(\chi+1)}$ (6.14) is not the same for long and short bursts’ data scattering on the $E_{p}$-$L_{\gamma}$ plane, as shown in Figure 7 (left) of the reference [25].

### v2

@@ 8 Collimated fireball intrinsic correlations
where the total mass and angular momentum ratio $J/M$ characterizes rotating systems in gravitational collapsing processes, the initial centre density $\rho^{\rm in}_{c}/\rho_{0}$ crucially relates to the gravitational binding energy gain, averaged thermal index $\gamma$ and material binding energy $\lambda$ represent material reactions in processes. The photon-pair collimated fireball energetic properties (6.8,6.9) and (6.13,6.15) explicitly depend on the initial and final core central densities $\rho_{c}^{\rm in}$ and $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}\approx 10\rho_{0}$. They determine the gravitational binding energy gain and conversion to baryon kinetic energy and photon-pair thermal energy in gravitational collapses. These characteristics should also be held for gravitational collapses of non-spherical cores or binary mergers by introducing effective initial and final central densities $(\rho_{c}^{\rm in})_{\rm eff}$ and $(\rho_{c}^{\rm max})_{\rm eff}$ of the system.
8.1 Intrinsic correlations among collimated fireball quantities
All collimated fireball quantities (6.8,6.9) and (6.13,6.14,6.15) are expressed as scaling laws of the initial core centre density $\rho_{c}^{\rm in}$ since it characterizes how much the gravitational binding energy gains when the systems of masses $M$ gravitationally collapse from $\rho_{c}^{\rm in}$ to final core centre density $\rho_{c}^{\rm fi}=\rho_{c}^{\rm max}\approx 10\rho_{0}$.
@@ 8.1 Intrinsic correlations among collimated fireball quantities
By simplifying out the initial core centre density $\rho_{c}^{\rm in}$ among Eqs. (6.8,6.9) and (6.13,6.15), we obtained theoretical correlations among the spectral peak $E_{p}$, total photon-pair energy $E_{\gamma}$ and peak luminosity $L_{\gamma}\propto cR_{\gamma}^{2}\langle T_{\gamma}\rangle^{4}\propto cE_{\gamma}/R_{\gamma}$ of the photon-pair sphere or collimation in a local rest frame,
@@
depending on the unique parameter $\gamma\gtrsim 1$ and $\chi=\frac{\gamma-1}{4\gamma-5}\sim{\mathcal{O}}(1)$  [^4] It is an inverse of the original $\chi-$definition [26]..
These collimated fireballs undergo ultra-relativistic expansion and interact with massive stellar material, forming ultra-relativistic ejecta whose Lorentz factors can be estimated as [10, 15, 17, 18, 19, 20, 21]
$\Gamma_{0}\approx E_{\gamma}/M_{\rm ej}$ (8.3)
where $M_{\rm ej}$ are ejected masses. This article does not present studies of ultra-relativistic outflow of collimated fireballs. However, we attempt to explain the ejected masses $M_{\rm ej}$ in this simplified model, thus to have an estimation of Lorentz factors (8.3).
At the end of homologous collapse, the homologous baryon core density $\rho$ (5.12) profile (Fig. 1) and photon-pair collimated fireball profile (Fig. 4) show that the baryon core size $r_{s}$ is larger than photon-pair collimated fireball size $r_{\gamma}$, e.g., $r_{s}>r_{\gamma}$ (6.6). This result implies that the photon-pair collimated fireball expansion erupts and pushes/engulfs the outer part (tail) of collapsing baryon cores outward. Therefore, the outer part of the baryon core density contributes to the ejecta masses $M_{\rm ej}\ll M$. We approximately assume the positive correlation between the ejecta mass $M_{\rm ej}$ and initial centre mass density $\rho^{\rm in}_{c}$, i.e., $M_{\rm ej}\propto(\rho^{\rm in}_{c})^{\beta}$. The reason is that the homologous core density profile shows that total collapsing core masses $M$ are large, centre mass densities $\rho^{\rm in}_{c}\propto M$ are large and outer part masses $M_{\rm ej}\propto M$ are large. The positive parameter $\beta$ is of order unity, and we set $\beta=1$ for simplicity.
The total collimated fireball energy is $E_{\gamma}\propto(\rho^{\rm in}_{c})^{(4\gamma-5)}$ (6.13). Therefore, the ejecta Lorentz factor is approximately given by
$\Gamma_{0}\approx E_{\gamma}/M_{\rm ej}\propto(\rho_{c}^{\rm in})^{(4\gamma-5)-1}.$ (8.4)
The collimated fireball has size scale $R_{\gamma}\propto(\rho_{c}^{\rm in})^{\frac{\gamma-2}{2}}$ (6.8). Its time scale $\tau_{\gamma}=R_{\gamma}/c$ relates to an observational one $T_{90}$ by the approximate relation $T_{90}\approx\tau_{\gamma}/(2\Gamma^{2}_{0})$ [44],
$T_{90}$$\approx$ $\tau_{\gamma}/(2\Gamma^{2}_{0})\propto(\rho_{c}^{\rm in})^{-\frac{2-\gamma}{2}-2(4\gamma-5)+2}.$ (8.5)
As a result, we approximately express the ejecta Lorentz factor $\Gamma_{0}$ and GRB observational time scale $T_{90}$ as scaling laws of the initial core centre density $\rho_{c}^{\rm in}$.
From the $\{\tau_{\gamma},E_{\gamma},\langle T_{\gamma}\rangle,L_{\gamma}\}$ universal $\chi$-correlations (Eqs. (7.2-7.7) in Ref. [26]), we obtain the $\{T_{90},E_{iso},E_{p},L_{iso}\}$ universal $\chi$-correlations:
$E_{p}$ $\propto$ $E_{iso}^{{}^{\chi}},$ (8.6)
$E_{p}$ $\propto$ $L_{iso}^{\frac{\chi}{3+\chi-\chi_{{}_{\Gamma}}}},$ (8.7)
$L_{iso}$ $\propto$ $E_{iso}^{3+\chi-\chi_{{}_{\Gamma}}}.$ (8.8)
and
$E_{p}$ $\propto$ $T_{90}^{-\frac{2\chi}{3(1+\chi)-2\chi_{{}_{\Gamma}}}},$ (8.9)
$E_{iso}$ $\propto$ $T_{90}^{-\frac{2}{3(1+\chi)-2\chi_{{}_{\Gamma}}}},$ (8.10)
$L_{iso}$ $\propto$ $T_{90}^{-2\frac{(3+\chi)-\chi_{\Gamma}}{3(1+\chi)-2\chi_{{}_{\Gamma}}}},$ (8.11)
where $\chi_{{}_{\Gamma}}=2/(4\gamma-5)$, and they can be recast as other correlations. Based on the simplified model, we obtain the theoretical correlations (8.6), (8.7) and (8.2), anti-correlations (8.9), (8.10) and (8.11) in a unified framework. They represent intrinsic properties of the photon-pair collimated fireballs and should be universal for all GRBs’ progenitors.
We cannot analytically obtain intrinsic correlations between the collimated fireball angle $\theta_{\rm coll}$ (7.2), the ratio $J/M$ and other collimated fireball quantities $E_{p}^{j}$, $E_{\gamma}^{j}$, $L_{\gamma}^{j}$ and $R_{\gamma}^{j}$. However, in the next Sec. 9, we will try to obtain the numerical trends (Table 1) of collimated fireball quantities in terms of rotation parameter $a^{2}\Omega^{2}_{c}\propto J/M$ at the end of homologous collapses. From these numerical trends, we infer collimated fireball intrinsic correlations, including the collimated fireball angle $\theta_{\rm coll}$. Henceforth, we omit the superscript j of collimated fireball quantities for simplifying notations unless otherwise specified.
8.2 Intrinsic correlations imprint on observed GRB data
The theoretically obtained intrinsic correlations (8.6) and (8.7) seem to give the Amati and Yonetoku empirical correlations in GRBs [45, 46], if we identify the photon-pair collimated fireball energy $E_{\gamma}$ and mean energy $\langle T_{\gamma}\rangle$ to the observed isotropic equivalent energy $E_{\rm iso}$ and peak energy $E_{p}$. However, we cannot simply make such identification for the reasons that a photon-pair collimated fireball is opaque and undergoes complex processes up to the transparency at which it becomes an observably relevant photosphere or collimated one. One of the reasons is that the peak energy $E_{p}$ value sensitively depends on the baryonic material loading (ejecta mass) $M_{\rm ej}$ on collimated fireballs, see for example, Refs. [18, 19, 20, 21]. However, we present a general discussion on the connections between theoretical correlations (8.6-8.11) and trends (Table 1) of different collimated fireballs and observational correlations of different GRB sources.
The collimated fireballs provide the energetic characteristics for observed GRB events. Suppose each collimated fireball created by a gravitational collapse corresponds to a GRB event/source observed. An initially opaque collimated fireball (initial state) undergoes complex (intermediate processes) to a transparent “photosphere” (final state) of GRB phenomena observed. The collimated fireball quantities $T_{\gamma}$, $E_{\gamma}$, $L_{\gamma}$ and $R_{\gamma}$ are not the same as the observed GRB event’s mean energy, total energy, total luminosity, and time duration. Therefore, the theoretical correlations (8.6-8.11) and trends (Table 1) cannot be examined by directly comparing them with the observed quantities in GRB events.
However, collimated fireball overall quantities ($T_{\gamma},E_{\gamma},L_{\gamma},R_{\gamma}$), intrinsic relations (8.6-8.11) and trends (Table 1) must imprint (code) themselves in observed GRB data. Although the intermediate processes from a collimated fireball to an observed GRB source are complex, they randomly differ from one GRB source to another. Such randomness implies that the observed GRB quantities, corresponding to collimated fireball counterparts, should randomly scatter around the intrinsic relations (8.6-8.11) and trends (Table 1) if these scaling relations and trends truly reflect the universal natures of centre engines for GRB energetics. Let’s take into account more data on GRB events. We will achieve a more statistically confident level on the validity of the intrinsic relations (8.6-8.11) and trends (Table 1) of collimated fireball quantities. These are only general discussions on the observational relevance of the intrinsic relations (8.6-8.11) and trends (Table 1) in the statistical sense by using a large sample of observed GRB data. For this purpose, the relevant GRB phase identification, spectrum and event selection, and elaborate data are required. These subjects are currently under investigation.
$a^{2}\Omega^{2}_{c}$ $\theta_{\rm coll}$ $E^{j}_{p}/E_{p}$ $E^{j}_{\gamma}/E_{\gamma}$ $L^{j}_{\gamma}/L_{\gamma}$ $R^{j}_{\gamma}/R_{\gamma}$ $N^{j}_{\gamma}/N_{\gamma}$ $V^{j}_{\gamma}/V_{\gamma}$
@@ 8.2 Intrinsic correlations imprint on observed GRB data
Table 1: For selected values of the rotation parameter $a^{2}\Omega^{2}_{c}$, which is proportionally related to the ratio $J/M$ (5.30), we tabulate the collimated angle $\theta_{\rm coll}$ (7.2) and the ratios of the spectral peak $E^{j}_{p}/E_{p}$, total energy $E^{j}_{\gamma}/E_{\gamma}$, maximal luminosity $L^{j}_{\gamma}/L_{\gamma}\propto(E^{j}_{\gamma}/E_{\gamma})(R_{\gamma}/R^{j}_{\gamma})$, minimal size $R^{j}_{\gamma}/R_{\gamma}$, total particle number $N^{j}_{\gamma}/N_{\gamma}$ and volume $V^{j}_{\gamma}/V_{\gamma}$ between the photon-pair spherical and collimated fireballs for comparisons. The collimated (with superscript $j$) and spherical fireball quantities are given by Eqs. (6.8,6.14,6.15). Their ratios do not explicitly depend on the initial (final) core centre density $\rho_{c}^{\rm in}$ ($\rho_{c}^{\rm fi}$) of gravitational collapses.
9 Long vs short GRBs by progenitor angular momentum
@@
9.1 Intrinsic trends of collimated fireball properties varying with $J/M$
Table 1 from left to right columns shows the trends of collimated fireball quantities varying as the angular momentum and mass ratio $J/M\propto a^{2}\Omega^{2}_{c}$ increases,
  1. (1)
    the collimated fireball angle $\theta_{\rm coll}$ becomes narrower;
  2. (2)
    the fireball energy spectrum $E^{j}_{p}$ becomes harder;
  3. (3)
    the fireball isotropic equivalent energy $E^{j}_{\gamma}$ becomes smaller;
  4. (4)
    the fireball isotropic equivalent luminosity $L^{j}_{\gamma}$ becomes larger;
  5. (5)
    the fireball characteristic size $R^{j}_{\gamma}$ (time $\tau^{j}_{\gamma}=R^{j}_{\gamma}/c$) becomes smaller (shorter).
We obtain these numerical results for massive slowly-rotating core collapses with the ratios $J/M\ll 1$.
Nevertheless, we extrapolate these trends in Table 1 to the larger $J/M$ values for fast-rotating cores and binary mergers, to gain an insight into the collimated fireballs generated in gravitational collapses of fast-rotating systems. The arguments are as follows:
@@ 9.1 Intrinsic trends of collimated fireball properties varying with $J/M$
    The microscopic physics processes of hadron collisions (Sec. 4) accounting for producing fireballs in slow- or fast-rotating massive collapsing cores and binary coalescence are the same. It occurs at nucleon interaction time and length scales when rotating massive cores and binary mergers run into the homologous collapsing phase around the nuclear density. The collimated fireball quantities are mainly determined by the initial centre density $\rho_{c}^{\rm in}$ of massive collapsing cores or the effective values $(\rho_{c}^{\rm in})_{\rm eff}$ of binary mergers. These microscopic processes at length scales of nuclear interactions play essential roles in fireballs’ formation.
@@
    The macroscopic physics processes of the dynamic interplay between gravitational binding and rotational repelling energy for collimating fireballs is the same for slow- or fast-rotating massive core collapses and binary mergers. The processes and results are characterized by the ratio $J/M$ of total angular momentum $J$ and mass $M$ of systems, which conserve in the entire core collapsing or binary merging process. Such overall ratio $J/M$ possesses the same physical meaning for a rotating core collapsing and binary merging into a rotating core collapsing near the nuclear density when fireballs occur. Therefore, the $J/M$-described macroscopic processes should play a similar key role in fireballs’ collimation for massive core collapses and binary mergers.
@@
    In addition, the obtained numerical trends of collimated fireball characteristic quantities vary monotonically in terms of $J/M$, following the expected physical behaviours from the dynamic interplay between gravitational binding and rotational repelling energy for collimating fireballs.
@@
    Numerical simulations show binary coalescence has a very different structure from core collapses. In binary coalescence cases, collimated fireballs should form in the most later gravitational collapsing phase of binary coalescence when binary stellar cores merge, becoming a fast-rotating core collapsing around the nuclear density. From these insights, we expect some similarity of fireballs’ collimation between core collapses and binary mergers in terms of their $J/M$. To demonstrate it, full numerical simulations in the GR framework are necessary.
However, these arguments are highly speculative rather than justified. We cannot conclude that the trends apply to fast-rotating collapsing systems and binary mergers relating to GRBs’ observations because of our analytical studies based on a simplified model of a slowly rotating and massive core collapsing in the Newtonian approximation. The key issue is how fireballs collimate in gravitation collapses of fast-rotating systems. It is very complicated to perform analytical studies on collimated fireballs formed in the gravitational collapses of fast-rotating massive core and binary systems in the GR framework. To obtain conclusively the properties of collimated fireballs varying in terms of $J/M$, one has to perform numerical studies in the GR framework to study fast-rotating collapsing systems and binary mergers.
Despite our lack of such numerical simulations, it is worthwhile to examine the self-consistency of these theoretically speculated $J/M$ trends (Table 1) of collimated fireballs from the empirical viewpoint of long and short GRBs’ observations.
9.2 Long vs short GRBs and jet angles
These collimated fireball quantities and trends in Table 1 cannot directly connect with observed quantities in GRBs data. However, as discussed in Sec. 8.2, these intrinsic trends (1)-(5) of energetic and geometric quantities varying from nearly spherical to collimated fireballs in terms of $J/M\propto a^{2}\Omega_{c}^{2}$ must imprint or code in the data of observed GRBs if these trends truly reflect the universal nature of collimated fireball energetics of GRB centre engines. In other words, the imprints of these trends on data should be significantly evident in a large sample of observed GRB data.
To see if the trends (1)-(5) and Table 1 are relevant to observations connecting with jet features, we need to assume the collimated fireball angle $\theta_{\rm coll}$ proportional relation (or positive correlation) (7.3) to the jet angle $\theta_{\rm jet}$ and observational one $\theta^{*}_{\rm jet}$, as discussed in Sec. 7.2. Based on the previous discussions on small $J/M$ for core collapses and large $J/M$ for binary mergers, we will examine the trends (1)-(5) in Table 1 as $J/M$ increase to see whether or not the trends and classifications by $J/M$ are consistent with observational data analysis of short vs long GRBs by a comprehensive study of the energetics and prompt gamma-ray correlations [47].
We observe that the trends (1)-(5) of collimated fireball quantities as increasing angular momentum and mass ratio $J/M$ in Table 1 support the common view that long GRBs from massive rotating core collapses (small $J/M$) and short GRBs from massive binary coalescence (large $J/M$). Based on the positive correlation $\theta_{\rm coll}\sim\theta_{\rm jet}\sim\theta^{*}_{\rm jet}$ (7.3), we find some novel features on observed ultra-relativistic jet angles $\theta^{*}_{\rm jet}$ correlating to long and short GRBs. The following qualitative observations and discussions are in order.
  1. (i)
    Due to the angular momentum $J/M$ and repulsive centrifugal potential increase, the gravitational binding energy gain becomes smaller. The ratios $E_{\gamma}^{j}/E_{\gamma}$ and $N_{\gamma}^{j}/N_{\gamma}$ are smaller than one but still in the order of unity, because most photons and pairs are produced at the centre ($z=0,y=0$) of the collimated fireball, see Figs. 4 and 5.
  2. (ii)
    The collimated fireball spectral peak $E_{p}^{j}/E_{p}$ significantly increases up to twice larger than the spherical one, because the photon-pair energy density increases as the collimated fireball volume $V^{j}$ decreases. As a result, the collimated fireball spectral peak $E^{j}_{p}$ increases, total energy $E^{j}_{\gamma}$ and characteristic time scale $R^{j}/c$ decrease as angular momentum $J/M$ increases. These trends might give a possible explanation for two distinct Amati correlations, relating to $E_{p}\propto E_{\gamma}^{\chi}$ (8.6), respectively for long and short bursts’ data scattering on the $E_{p}$-$E_{\gamma}$ plane, see Fig 7 (right) of the reference [47], where two distinct Amati relations are observed for long and short bursts.
  3. (iii)
    The collimated fireball spectral peak $E^{j}_{p}$ and the peak luminosity $L^{j}_{\gamma}$ increase as angular momentum $J/M$ increases. It implies the Yonetoku correlation, relating to $E_{p}\propto L_{\gamma}^{\chi/(\chi+1)}$ (8.2), is not the same for long and short bursts’ data scattering on the $E_{p}$-$L_{\gamma}$ plane, as shown in Figure 7 (left) of the reference [47].
  4. (iv)